OpenStudy (anonymous):

What is the solution of log(2x - 3)125 = 3 ?

OpenStudy (lgbasallote):

first change it into exponential form \[(2x - 3)^3 = 125\] then get the cube root of both sides \[\sqrt[3]{(2x-3)^3} = \sqrt[3]{125}\] \[2x - 3 = 5\] now it's easy to solve for x

OpenStudy (lgbasallote):

do you get that?

OpenStudy (anonymous):

Yeah, is the answer 4?

OpenStudy (lgbasallote):

yup